OpenStudy (anonymous):

What is the value of the expression when a = 5 and b = 3? 4a^2 - 5b^2

OpenStudy (anonymous):

Substitute a=5 and b=3 into your equation like so: 4*(5)^2 - 5*(3)^2 = 4*25 - 5*9 = 100-45 = 55

OpenStudy (anonymous):

oh , i figured because at first i was doing 4 x a then i was getting an wrong answer , THANKS !

OpenStudy (anonymous):

yeah its a little tricky because you have to square it, its easy to just rewrite the letter as an empty bracket with the squared symbol, then put your number in to avoid accidentally not squaring :)

OpenStudy (anonymous):

yea , ah thanks tho

OpenStudy (anonymous):

can u help me with this one | What is the value of the expression 3 ^–4?

OpenStudy (anonymous):

when you see a negative sign in the index (the power) it means you have to invert the expression. So the expression becomes 1/(3^4), because when you invert the negative disappears. Simplifying this it now becomes 1/81, which is approx. 0.0123456789...

OpenStudy (anonymous):

that is actually a really awesome number haha

OpenStudy (anonymous):

*also by invert i mean take the reciprocal

OpenStudy (anonymous):

yea i got that lmao yea it is awesome number

OpenStudy (anonymous):

the answer choices is | -81 | -12 | -1/81 | 1/81 tho

OpenStudy (anonymous):

choose 1/81

OpenStudy (anonymous):

the negatives are there to trick you

OpenStudy (anonymous):

a negative index does not make a negative number

OpenStudy (anonymous):

yea i figured , i just got 1/81 to haha stupid negatives

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

are you actually an substitute because them are kinda like teacher answers and teacher help

OpenStudy (anonymous):

nah im just a high level maths student haha

OpenStudy (anonymous):

ok i got this question 1/4^-3 and i got .015625 out of 4^-3 and divide it with 1 and got 64 . is that right ?

OpenStudy (anonymous):

yeah that's correct, good job

OpenStudy (anonymous):

:) i learned it from @substitute lol

OpenStudy (anonymous):

haha well what a great teacher you had :P

OpenStudy (anonymous):

ok last one i need help with please | What value of p solves the equation? p^-5 = 1/243

OpenStudy (anonymous):

ok so on the 'p' side of the equation, you want to invert it so you get 1/p^5 = 1/243. Do you see how then p^5=243? well then you just take the \[\sqrt[5]{243}\] and that's just 3

OpenStudy (anonymous):

ok i got another one like that its | 3^n = 1/81 and i got 27 , is that correct ?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

you need to convert 1/81 to a power of three so the indexes of the expressions on either side of the equations are equal

OpenStudy (anonymous):

1/81=1/3^4=3^-4 therefore if 3^n=3^-4 then n=-4

OpenStudy (anonymous):

ooh , god darn you good , Thank so much that was all the problems i needed help with | Go @substitute :")

OpenStudy (anonymous):

i think you might be better at math and explaning then my current math teacher lol

OpenStudy (anonymous):

haha no problem :) that wouldn't be a first, math teachers can sometimes be horrible

OpenStudy (anonymous):

yes ,especially in georgia where i live

OpenStudy (anonymous):

Hey i will have some more problems probably for you tomorrow because its 6:05am where i am and i havent been to sleep all day and im getting off os cya